What is the slope of a line perpendicular to the line 4x - 3y = 5?
Answer by mananth(16946)

4 x -3 y = 5
Find the slope of this line

-3 y = -4 x + 5
Divide by -3
y = 4/3 x -1.67
Compare this equation with y=mx+b
slope m = 4/3
m1*m2=-1
The slope of a line perpendicular to the above line will be the negative reciprocal -3/4